OpenStudy (anonymous):

+2 - -19 = ? Need help!!!!

OpenStudy (bahrom7893):

+2 + 19 = 21

OpenStudy (bahrom7893):

double negative is a positive

Parth (parthkohli):

Because -(-blah) = +blah,\[+2-(-19) = +2 + 19\]

OpenStudy (anonymous):

\[12-(-19)=12+19= ?\]

OpenStudy (anonymous):

So what's the answer

OpenStudy (anonymous):

31 :)

OpenStudy (anonymous):

I think it's +21

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

Ohh. Yeap .. I calculated it according to my post.. Sorry I have written 12 instead of 2.. :D

OpenStudy (anonymous):

Ok that you

OpenStudy (anonymous):

+16 - -34 = ?

OpenStudy (anonymous):

Same as the above.. Give a try.. :)

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

Nope. 16+34=?

OpenStudy (anonymous):

@kiki41554

OpenStudy (anonymous):

+50

OpenStudy (anonymous):

Correct. Well done :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You are always welcome... :)

OpenStudy (anonymous):

-11 - -47= ??

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

- - = + Try to do it...

OpenStudy (anonymous):

So I add them

OpenStudy (anonymous):

But the minus of the first one wouldn't change.. - a - -b = -a+b

OpenStudy (anonymous):

Yeah add them up. Don't forget the minus mark at the front..